# The currently observed clumps cannot be the ”direct” precursors of the currently observed open clusters

### v1

@@ 3.3.2 Timescale of gas free
Seven of the 17 MSFRs investigated in the MYStIX project are covered in the ATLASGAL survey (Schuller et al., 2009), which are the Lagoon nebula, NGC 6334, NGC 6357, the Eagle nebula, M 17, the Carina nebula, and the Trifid nebula. As shown in Sec.A, the ATLASGAL+Planck or only ATLASGAL 870 m data (Csengeri et al., 2016) were used to trace the gas distribution around embedded or very young clusters. On these maps, clusters and clumps are well separated. However, NGC 6334 and M 17 still have abundant gas, which means they are younger than other MSFRs. As shown in Fig.3, the clusters in NGC 6334 and M 17 (”a”) indeed have smaller ages (the median value 1.4 Myr) than those in other MSFRs (”b”, the median value 2.2 Myr). If embedded clusters in NGC 6334 and M 17 just finished the gas expulsion, then this implies an upper limit on gas expulsion time around 1.4 Myr.
@@ 3.3.3 Slow and moderate gas expulsions
Embedded clusters form in clumps. More massive clumps can produce more massive clusters, leading to stronger feedback and larger gas expulsion velocity (Dib et al., 2013). There should be correlations between the feedback strength, the clump (or cluster) mass and the gas expulsion velocity (). And low-mass clusters should have a slower gas expulsion process compared with high-mass clusters. As shown in Pang et al. (2021), low-mass clusters tend to agree with the simulations of slow gas expulsion. Except for the feedback strength, the SFE determines the total amount of the remaining gas, which also influences the timescale of gas expulsion. Since the gas expulsion process is driven by feedback, and the effectiveness of feedback will depend on the geometric shape of the gas shell surrounding the embedded cluster (Wünsch et al., 2010; Rahner et al., 2017). Therefore, a complex or non-spherical gas distribution is also a factor to change the timescale of gas expulsion. In short, these parameter uncertainties can ultimately be incorporated into the timescale of gas expulsion. Thus, apart from the fast gas expulsion with the gas decay time described above, we also simulated slow and moderate gas expulsions. As described in Sec.3.3.2, there may be an upper limit on the gas expulsion timescale, 1.4 Myr. Therefore, for the slow gas expulsion, the gas decay time is set to be 10. The moderate gas expulsion is between the fast and slow gas expulsions. Considering the large parameter space between and 10, we simulated two kinds of moderate gas expulsions, i.e. 2 (”moderate2”) and 5 (”moderate1”) . Fig.4 shows the evolution of gas mass under different gas expulsion modes over time.
@@ 3.3.4 Star formation efficiency
The total amount of the residual gas not only affects the gas expulsion timescale, but it also significantly influences the strength of the gas potential. The strength of the external gas potential may have a considerable impact on the evolution of embedded star clusters. The total amount of the residual gas is determined by the SFE of the clump. In the simulations above, we assume SFE = 0.33 for all star clusters. In Zhou et al. (2024), we calculated the SFE of ATLASGAL clumps with HII regions, and found a strong anti-correlation between the SFE and the clump mass. Now, we directly derive the residual gas mass using equation.6 in Zhou et al. (2024), which encodes the variation of SFE with clump mass, i.e.
@@ 3.3.5 Procedure
The McLuster program (Küpper et al., 2011) was used to set the initial configurations. The dynamical evolution of the model clusters was computed using the state-of-the-art PeTar code (Wang et al., 2020). PeTar employs well-tested analytical single and binary stellar evolution recipes (SSE/BSE) (Hurley et al., 2000, 2002; Giacobbo et al., 2018; Tanikawa et al., 2020; Banerjee et al., 2020). In PeTar, there are some important parameters, i.e. the tree time step (-s), the changeover radius (-r), the neighbor searching radius (–r-search-min) and the multiple group radius (–r-bin). The tree time step is a fixed time step to calculate the long-range (particle-tree) force. The changeover region is the overlap shell between the long-range and short-range interaction. Especially, -s and -r significantly affect the accuracy and performance of the simulations. Without specifying -s and -r, PeTar determines them automatically assuming the input model is a sphercial symmetric star cluster with a King or a Plummer like density profile. The neighbor searching radius is defined for each particle to identify neighboring candidates expected to be within the changeover region during the subsequent tree time step. Upon determining the changeover radius (-r), PeTar automatically computes the neighbor searching radius. The multiple group radius (–r-bin) is used as the criterion for selecting group members, which is automatically determined based on the changeover inner radius. In this work, all these parameters are automatically determined by PeTar.
@@ 3.4 Simulation results analysis
In Fig.5, as expected, the mass loss of the clusters strongly depends on the gas expulsion modes. In the scenario of fast gas expulsion, gas is removed almost instantaneously, leading to a strong impact on the clusters. In contrast, with slow gas expulsion, the gas disperses over a timescale significantly longer than the cluster’s crossing (dynamical) time, enabling the cluster to adapt to the changing gravitational potential and expand while maintaining near dynamical equilibrium, as discussed in Baumgardt & Kroupa (2007). This quasi-static (or adiabatic) evolution results in substantially less stellar mass loss, as presented in Fig.5. Therefore, it best fits the mass distribution of the observed open clusters. However, it fails to match the radius distribution, as shown in Fig.6(d1). For the fast gas expulsion, the clusters can not survive more than 100 Myr. Similarly, in Kroupa (2001); Brinkmann et al. (2017), the massive cluster containing many O stars most likely survives as an open cluster. Low-mass embedded clusters or groups dissolve quickly, partly owing to loss of their residual gas. Lower SFE means more residual gas, leading to more intense expansion and mass loss of the cluster after gas expulsion, as shown in Fig.5(b). Actually, a lower SFE is equivalent to a shorter gas expulsion timescale.
For high-mass clusters, such as 3000 M⊙, and 2 will give significantly different results. Thus, the simulations are very sensitive to the gas decay time. We need to carefully constrain this parameter from observation in future work. Since with being constrained by equation.1, we can focus on the measurement of the gas expulsion velocity () in observation, especially the correlation between and the clump (or cluster) mass, as discussed in Sec.3.3.3.
In Fig.6, in order to explain the radius distribution of the observed open clusters, fast gas expulsion is required, but to account for the mass distribution, slow gas expulsion is necessary. This contradiction makes both fast and slow gas expulsions inappropriate. To reconcile this contradiction, the gas expulsion should be somewhere between fast and slow. In terms of both mass and radius, ”moderate2” can best fit the observed open clusters, and the main contribution is from the high-mass cluster (3000 M⊙). To explain the mass and radius distributions of the observed open clusters, initial embedded clusters with masses more than 3000 M⊙ are necessary. However, in Fig.1, the upper limit of the embedded cluster sample is less than 1000 M⊙. And only few clumps have a mass more than 3000 M⊙. For an open cluster, its progenitor should have a significantly larger mass. Thus, the currently observed clumps cannot be the ”direct” precursors of the currently observed open clusters.
@@ 4 Discussion
There may be two evolutionary tracks from embedded clusters to open clusters, i.e. single embedded cluster evolution scenario and multiple embedded clusters merger scenario. In Zhou et al. (2024b), we used the minimum spanning tree method to measure the separations between clusters and gas clumps in each MSFR shown in Sec.A. The separations between clusters, between clumps, and between clusters and clumps are comparable, which indicates that the evolution from clump to embedded cluster proceeds in isolation and locally, and does not affect the surrounding objects significantly. Using Gaia DR2 data, Kuhn et al. (2019) investigated the kinematics of subclusters identified in Kuhn et al. (2014) for the same MSFRs, and found no evidence that these groups are merging. Therefore, in this work, we only focus on the first scenario. However, as presented above, the masses of current clumps or embedded clusters in the Milky Way are insufficient to evolve into the observed open clusters.
@@
For the analysis in this work, we assumed a constant star formation history (SFH) of the Milky Way. If there is a burst-like and time-dependent SFH of the Milky Way, the currently observed open clusters may be remnants of massive clumps from the past. The epoch of peak star formation rate density is thought to lie between redshifts of 2 and 3 (Madau et al., 1998; Hopkins & Beacom, 2006; Bouwens et al., 2011; Moster et al., 2013). Gaia DR2 reveals a star formation burst in the Galactic disc 2-3 Gyr ago (Mor et al., 2019). As an inference, the peak period of star formation in the Milky Way has passed, and most of the massive clumps have evolved into open clusters or older clusters. At present, only relatively low-mass clumps remain. This possibility is worth exploring further in future work.
@@
Another explanation is that those massive and relatively old open clusters come from multiple embedded clusters mergers. As shown in Sec.A, those embedded clusters or very young clusters in each MSFR are very near. There is a great possibility that they will merge after expansion. One molecular cloud can produce multiple embedded clusters, which may merge into larger clusters. This scenario is consistent with the finding in Sec.3.2, i.e. most molecular clouds may only form one open cluster. In a forthcoming work (Zhou et. al. in preparation), we will simulate the merger between subclusters in each MSFR, with the initial conditions of the simulations directly derived from the observation. Actually, there is now extensive literature arguing that star clusters form from mergers between embedded clusters, both from simulations (Fujii et al., 2012; Howard et al., 2018; Sills et al., 2018; Fujii et al., 2022; Dobbs et al., 2022; Cournoyer-Cloutier et al., 2023; Reina-Campos et al., 2024) and observations (Sabbi et al., 2012; Pang et al., 2022; Della Croce et al., 2023a; Fahrion & De Marchi, 2024). A comparison with these works will also be presented in the forthcoming work.
@@
In this work, the initial density profile of all clusters is the Plummer profile. However, observations of embedded clusters show clusters with high ellipticities (Kuhn et al., 2014; Getman et al., 2018). Due to the limitations of the method used here, we are currently unable to discuss the impact of the cluster’s ellipticity on its subsequent evolution. For now, we have only performed N-body simulations. In future work, it is necessary to compare these with more comprehensive hydrodynamics+N-body simulations, such as Fujii et al. (2021); Rieder et al. (2021); Polak et al. (2024), to further test and constrain the initial conditions used in the N-body simulations presented in this work.
@@ 5 Conclusion
In this work, we collected samples of Galactic clumps, embedded clusters, and open clusters to compare their physical properties. Overall, the radii of open clusters are significantly larger than those of embedded clusters and clumps. The radii of embedded clusters are larger than those of clumps, which may be due to the expansion of embedded clusters. The mass distribution of open clusters covers a significantly larger mass range than that of embedded clusters, by about one order of magnitude. Given the current mass distribution of clumps in the Milky Way, it is unlikely that a single clump evolving into a single embedded cluster, and subsequently into a single open cluster, could account for the observed open clusters with old ages and large masses, which is also supported by N-body simulations of individual embedded clusters.
We used standard N-body simulations as in previous works to interpret the physical parameters of the observed open clusters. The mass loss of the simulated clusters strongly depends on the gas expulsion modes. For the fast gas expulsion, the clusters (3000 M⊙) can not survive more than 100 Myr. In order to explain the radius distribution of the observed open clusters, fast gas expulsion is required, but to account for the mass distribution, slow gas expulsion is necessary. This contradiction makes both fast and slow gas expulsions inappropriate. To reconcile this contradiction, the gas expulsion should be somewhere between fast and slow. Anyway, to explain the mass and radius distributions of the observed open clusters, initial embedded clusters with masses more than 3000 M⊙ are necessary. However, in Fig.1, the upper limit of the embedded cluster sample is less than 1000 M⊙. And only few clumps have a mass larger than 3000 M⊙. For an open cluster, its progenitor should have a significantly larger mass. Thus, the currently observed clumps cannot be the ”direct” precursors of the currently observed open clusters.
If there is a burst-like and time-dependent star formation history of the Milky Way, the currently observed open clusters may be remnants of massive clumps from the past. At present, only relatively low-mass clumps remain. Another explanation is that those massive and relatively old open clusters come from multiple embedded clusters mergers. We compared the separation of open clusters and the typical size of molecular clouds in the Milky Way, and found that most molecular clouds may only form one open cluster. Each molecular cloud typically includes many embedded clusters, the typical separation between them can be 1 pc (Zhou et al., 2024b). Thus, after expansion, they should be able to undergo mergers. Further study is necessary to distinguish between different scenarios.
@@ Acknowledgements.
Thanks the referee for providing detailed and constructive review comments, which have helped to improve and clarify this work. We would like to thank Pavel Kroupa for the comments and suggestions that helped improve and clarify this work.References

### v2

@@ 3.3.2 Timescale of gas free
Seven of the 17 MSFRs investigated in the MYStIX project are covered in the ATLASGAL survey (Schuller et al., 2009): the Lagoon nebula, NGC 6334, NGC 6357, the Eagle nebula, M 17, the Carina nebula, and the Trifid nebula. As shown in Appendix A, the ATLASGAL+Planck or only ATLASGAL 870 m data (Csengeri et al., 2016) were used to trace the gas distribution around embedded or very young clusters. On these maps, clusters and clumps are well separated. However, NGC 6334 and M 17 still have abundant gas, which means they are younger than the other MSFRs. As shown in Fig. 3, the clusters in NGC 6334 and M 17 (”a”) are indeed younger (the median value is 1.4 Myr) than those in other MSFRs (”b”; the median value is 2.2 Myr). If embedded clusters in NGC 6334 and M 17 just finished the gas expulsion, then this implies an upper limit on the gas expulsion time of around 1.4 Myr.
@@ 3.3.3 Slow and moderate gas expulsions
Embedded clusters form in clumps. More massive clumps can produce more massive clusters, leading to stronger feedback and a higher gas expulsion velocity (Dib et al., 2013). There should be correlations between the feedback strength, the clump (or cluster) mass, and the gas expulsion velocity (). And low-mass clusters should have a slower gas expulsion process compared with high-mass clusters. As shown in Pang et al. (2021), low-mass clusters tend to agree with the simulations of slow gas expulsion. Except for the feedback strength, the SFE determines the total amount of the remaining gas, which also influences the timescale of gas expulsion. The gas expulsion process is driven by feedback, and the effectiveness of the feedback will depend on the geometric shape of the gas shell surrounding the embedded cluster (Wünsch et al., 2010; Rahner et al., 2017). Therefore, a complex or nonspherical gas distribution would also change the timescale of gas expulsion. In short, these parameter uncertainties can ultimately be incorporated into the timescale of gas expulsion. Thus, apart from the fast gas expulsion with the gas decay time, , described above, we also simulated slow and moderate gas expulsions. As described in Sect. 3.3.2, the upper limit on the gas expulsion timescale may be 1.4 Myr. Therefore, for the slow gas expulsion, the gas decay time was set to 10. The moderate gas expulsion is between the fast and slow gas expulsions. Considering the large parameter space between and 10, we simulated two kinds of moderate gas expulsions: 2 (”moderate2”) and 5 (”moderate1”) . Figure 4 shows the evolution of gas mass under different gas expulsion modes over time.
@@ 3.3.4 Star formation efficiency
The total amount of the residual gas not only affects the gas expulsion timescale, but it also significantly influences the strength of the gas potential. The strength of the external gas potential may have a considerable impact on the evolution of embedded star clusters. The total amount of residual gas is determined by the SFE of the clump. In the simulations described above, we assumed a SFE = 0.33 for all star clusters. In Zhou et al. (2024), we calculated the SFE of ATLASGAL clumps with HII regions and found a strong anticorrelation between the SFE and the clump mass. Here, we directly derived the residual gas mass using Eq. 6 in Zhou et al. (2024), which encodes the variation in SFE with clump mass:
@@ 3.3.5 Procedure
The McLuster program (Küpper et al., 2011) was used to set the initial configurations. The dynamical evolution of the model clusters was computed using the state-of-the-art PeTar code (Wang et al., 2020). PeTar employs well-tested analytical single and binary stellar evolution recipes (SSE/BSE) (Hurley et al., 2000, 2002; Giacobbo et al., 2018; Tanikawa et al., 2020; Banerjee et al., 2020). In PeTar, there are some important parameters: the tree time step (-s), the changeover radius (-r), the neighbor searching radius (–r-search-min), and the multiple group radius (–r-bin). The tree time step is a fixed time step used to calculate the long-range (particle-tree) force. The changeover region is the overlap shell between the long-range and short-range interaction. In particular, -s and -r significantly affect the accuracy and performance of the simulations. If -s and -r are not specified, PeTar determines them automatically, assuming that the input model is a spherical symmetric star cluster with a King- or a Plummer-like density profile. The neighbor searching radius is defined for each particle to identify neighboring candidates expected to be within the changeover region during the subsequent tree time step. Upon determining the changeover radius (-r), PeTar automatically computes the neighbor searching radius. The multiple group radius (–r-bin) is used as the criterion for selecting group members, which is automatically determined based on the changeover inner radius. In this work, all these parameters were automatically determined by PeTar (see Table LABEL:para).
@@ 3.4 Simulation results analysis
In Fig. 5, as expected, the mass loss of the clusters strongly depends on the gas expulsion modes. In the scenario of fast gas expulsion, gas is removed almost instantaneously, which strongly impacts the clusters. In contrast, with slow gas expulsion, the gas disperses over a timescale significantly longer than the cluster’s crossing (dynamical) time, enabling the cluster to adapt to the changing gravitational potential and expand while maintaining near dynamical equilibrium, as discussed in Baumgardt & Kroupa (2007). This quasi-static (or adiabatic) evolution results in substantially less stellar mass loss, as shown in Fig. 5. Therefore, it best fits the mass distribution of the observed OCs. However, it fails to match the radius distribution, as shown in Fig. 6(d1). For the fast gas expulsion, the clusters cannot survive more than 100 Myr. Similarly, in Kroupa (2001) and Brinkmann et al. (2017), the massive cluster containing many O stars most likely survives as an OC. Low-mass embedded clusters or groups dissolve quickly, in part due to the loss of their residual gas. A lower SFE means more residual gas, leading to more intense expansion and mass loss of the cluster after gas expulsion, as shown in Fig. 5(b). A lower SFE is equivalent to a shorter gas expulsion timescale.
For high-mass clusters, such as 3000 M⊙, and 2 will give significantly different results. Thus, the simulations are very sensitive to the gas decay time. We need to carefully constrain this parameter from observations in future work. Since , with being constrained by Eq. 1, we can focus on the measurement of the gas expulsion velocity () in observations, especially the correlation between and the clump (or cluster) mass, as discussed in Sect. 3.3.3.
In Fig. 6, fast gas expulsion is required to explain the radius distribution of the observed OCs, but to account for the mass distribution, slow gas expulsion is necessary. This contradiction makes both fast and slow gas expulsions inappropriate. To reconcile this contradiction, the gas expulsion should be somewhere between fast and slow. In terms of both mass and radius, ”moderate2” can best fit the observed OCs, and the main contribution is from the high-mass cluster (3000 M⊙). To explain the mass and radius distributions of the observed OCs, initial embedded clusters with masses higher than 3000 M⊙ are necessary. However, in Fig. 1, the upper limit of the embedded cluster sample is less than 1000 M⊙, and only a few clumps have a mass higher than 3000 M⊙. For an OC, its progenitor should have a significantly higher mass. Thus, the currently observed clumps cannot be the ”direct” precursors of the currently observed OCs.
@@ 4 Discussion
There may be two evolutionary tracks from embedded clusters to OCs: a single embedded cluster evolution scenario and a multiple embedded clusters merger (post-gas expulsion coalescence) scenario. In Zhou et al. (2024b), we used the MST method to measure the separations between clusters and gas clumps in each MSFR shown in Appendix A. The separations between clusters, between clumps, and between clusters and clumps are comparable, which indicates that the evolution from clump to embedded cluster proceeds in isolation and locally, and does not significantly affect the surrounding objects. Using Gaia DR2 data, Kuhn et al. (2019) investigated the kinematics of subclusters identified in Kuhn et al. (2014) for the same MSFRs and found no evidence that these groups are merging. Therefore, in this work, we only focused on the first scenario. However, as presented above, the masses of current clumps or embedded clusters in the Milky Way are too low to have evolved into the observed OCs.
@@
For the analysis in this work, we assumed a constant star formation history for the Milky Way. If its star formation history is instead burst-like and time-dependent, the currently observed OCs may be remnants of massive clumps from the past. The epoch of peak star formation rate density is thought to lie between redshifts of 2 and 3 (Madau et al., 1998; Hopkins & Beacom, 2006; Bouwens et al., 2011; Moster et al., 2013). Gaia DR2 reveals a star formation burst in the Galactic disk 2-3 Gyr ago (Mor et al., 2019). As an inference, the peak period of star formation in the Milky Way has passed, and most of the massive clumps have evolved into OCs or older clusters. At present, only relatively low-mass clumps remain. This possibility is worth exploring further in future work.
@@
Another explanation is that these massive and relatively old OCs come from post-gas expulsion coalescence of multiple embedded clusters. As shown in Appendix A, the embedded clusters or the very young clusters are very near to each other in each MSFR. There is a very real possibility that they will merge after expansion. One molecular cloud can produce multiple embedded clusters, which can merge into larger clusters. This scenario is consistent with the finding in Sect. 3.2, that most molecular clouds can only form one OC. In a forthcoming work (Zhou et. al. in preparation), we will simulate the post-gas expulsion coalescence of subclusters in each MSFR, with the initial conditions of the simulations directly derived from the observation. There is now extensive literature arguing that star clusters form from mergers between embedded clusters, both from simulations (Fujii et al., 2012; Howard et al., 2018; Sills et al., 2018; Fujii et al., 2022; Dobbs et al., 2022; Cournoyer-Cloutier et al., 2023; Reina-Campos et al., 2024) and observations (Sabbi et al., 2012; Pang et al., 2022; Della Croce et al., 2023a; Fahrion & De Marchi, 2024). A comparison with these works will also be presented in the forthcoming work.
@@
In the current work, the initial density profile of all clusters is the Plummer profile. However, observations of embedded clusters show clusters with high ellipticities (Kuhn et al., 2014; Getman et al., 2018). Due to the limitations of the method used here, we are currently unable to discuss the impact of the cluster’s ellipticity on its subsequent evolution. For now, we have only performed N-body simulations. In future work, it is necessary to compare them with more comprehensive hydrodynamics+N-body simulations, such as those presented in Fujii et al. (2021), Rieder et al. (2021), and Polak et al. (2024), to further test and constrain the initial conditions used in the N-body simulations presented in this work.
@@ 5 Conclusion
We collected samples of Galactic clumps, embedded clusters, and OCs to compare their physical properties. Overall, the radii of OCs are significantly larger than those of embedded clusters and clumps. The radii of embedded clusters are larger than those of clumps, which may be due to the expansion of embedded clusters. The mass distribution of OCs covers a significantly larger mass range than that of embedded clusters, by about one order of magnitude. Given the current mass distribution of clumps in the Milky Way, it is unlikely that a single clump evolving into a single embedded cluster, and subsequently into a single OC, could account for the observed OCs with old ages and high masses, which is also supported by N-body simulations of individual embedded clusters.
We used standard N-body simulations as in previous works to interpret the physical parameters of the observed OCs. The mass loss of the simulated clusters strongly depends on the gas expulsion modes. For the fast gas expulsion, the clusters (3000 M⊙) cannot survive more than 100 Myr. Fast gas expulsion is required to explain the radius distribution of the observed OCs, but to account for the mass distribution, slow gas expulsion is necessary; the real gas expulsion speed should be somewhere in between. In any case, to explain the mass and radius distributions of the observed OCs, initial embedded clusters with masses higher than 3000 M⊙ are necessary. However, in Fig. 1, the upper limit of the embedded cluster sample is less than 1000 M⊙, and only a few clumps have a mass higher than 3000 M⊙. The progenitor of an OC should have a significantly higher mass. Thus, the currently observed clumps cannot be the ”direct” precursors of the currently observed OCs.
If the star formation history of the Milky Way is burst-like and time-dependent, the currently observed OCs may be remnants of massive clumps from the past. At present, only relatively low-mass clumps remain. Another explanation is that those massive and relatively old OCs come from post-gas expulsion coalescence of multiple embedded clusters. We compared the separation of OCs and the typical size of molecular clouds in the Milky Way and find that most molecular clouds can only form one OC. Each molecular cloud typically includes many embedded clusters; the typical separation between them is 1 pc (Zhou et al., 2024b). Thus, after expansion, they should be able to undergo mergers. Further study is necessary to distinguish between the different scenarios.
@@ Acknowledgements.
Thanks the referee for providing detailed and constructive review comments, which have helped to improve and clarify this work.References